5x - 15y + z = 22
-10x + 12y – 2z = -8
4x – 2y – 3z = 9
Help me solve this problem

Answer by mananth(16946)

5x-15y+ z=22 -------------- 1
-10x+12y-2z=-8 -------------- 2
4x-2y-3z=9 -------------- 3

consider equation 1 &2 Eliminate y
Multiply 1 by 4
Multiply 2 by 5
we get
20x-60y +4z = 88
-50x + 60y-10z = -40
Add the two
-30x-6z = 48 -------------4
consider equation 2 & 3 Eliminate y
Multiply 2 by 1
Multiply 3 by 6
we get
-10x+12y-2z = -8
24x-12y -18z = 0
Add the two
14x -20z = -8 -------------5 5
Consider (4) & (5) Eliminate x
Multiply 4 by 14
Multiply (5) by 30
we get
-420x -84z = 672
420x-600z=-240
Add the two
-684z = 432
/ -684
z = -0.63 OR - 12/19

Plug the value of z in 5
14x + 12.63 = -8
14x = -20.63
x = -1.47 OR -1 9/19
plug value of x & z in 1
-7.37 + -15y + -0.63 = 22
-15y = 7.37 + 0.63 + 22
y= -2 OR -2

CHECK Equation 1
5 -1.47 -15 -2 1 -0.63
-7.37 30 -0.63 22